# Studying the Effects of Collaboration in Interactive Theme Discovery Systems

### v1

@@ Abstract
NLP-assisted solutions have gained considerable traction to support qualitative data analysis. However, there does not exist a unified evaluation framework that can account for the many different settings in which qualitative researchers may employ them. In this paper, we take a first step in this direction by proposing an evaluation framework to study the way in which different tools may result in different outcomes depending on the collaboration strategy employed. Specifically, we study the impact of synchronous vs. asynchronous collaboration using two different NLP-assisted qualitative research tools and present a comprehensive analysis of significant differences in the consistency, cohesiveness, and correctness of their outputs.
@@ 1 Introduction
Making sense of large amounts of textual data is a common task in many academic disciplines and is generally achieved through qualitative methods such as Thematic Analysis (Braun and Clarke, 2006) and Grounded Theory (Glaser et al., 1968). In these methods, researchers manually search through documents to identify common abstract themes that emerge from the data in a process commonly referred to as "inductive coding". However, manual inductive coding becomes untenable as datasets increase exponentially in size, leading researchers to adopt Natural Language Processing (NLP) solutions to automate parts of the coding process Brady (2019); Hilbert et al. (2019).
@@
In the last decade, several NLP solutions have been proposed to support qualitative research methods. These systems offer different ways to automatically wrangle large amounts of textual data to identify common patterns or themes, such as discovering hidden semantic structures using topic models Smith et al. (2018); Fang et al. (2023), grouping documents based on similarity and broadcasting small sets of human annotations to the entire dataset Pacheco et al. (2023), or providing live recommendations to users as they manually code the data Dai et al. (2023); Gao et al. (2024). To ensure the agency of the researchers, all of these systems seek to strike a balance between automation and manual coding by incorporating some type of human-in-the-loop (HitL) strategy.
@@
In most cases, these tools are evaluated to highlight the strengths and weaknesses of the proposed system. Examples include evaluating topic coherence with and without human input Fang et al. (2023) and contrasting machine-assisted code-book generation with a manual-only code-book generation Dai et al. (2023). In this paper, we argue that these evaluation strategies are often too narrow and do not account for the different real-world settings in which qualitative researchers employ HitL systems. In particular, we want to study how tools for inductive coding may result in different outcomes in different collaborative settings.
@@
Collaborative data analysis occurs when two or more researchers pool their focus into a shared body of data to produce an agreed interpretation. Potential advantages of collaboration include producing more reliable results, a richer understanding of the data, and knowledge that transcends individual perspectives Flick (2014). While there are many ways in which collaboration can occur, in this work we look at two common but opposite qualitative research collaboration settings. The first setting is defined by individual coders working separately before jointly consolidating their findings (asynchronous coding). The second setting is defined by coders working in groups and identifying themes through synchronous discussion (synchronous coding). Each setting has different strengths: asynchronous coding can be a more flexible option for teams as they can work across different regions and time zones, whereas synchronous coding can combine diverse perspectives and streamline the annotation processes.
@@
We test these settings by choosing two NLP tools for inductive coding, and performing coding experiments in the synchronous and asynchronous contexts. We choose tools that use different underlying NLP methods for this work: (1) the Fang et al. (2023) system, which uses a traditional human-in-the-loop topic modeling solution, and (2) the Pacheco et al. (2023) system, which models emerging themes with respect to their relationship to other high-level concepts, groups similar instances based on this, and leaves theme definitions and annotations to users, which are then propagated to the full dataset. This way, we can have a more representative understanding of how different tools can be affected in synchronous and asynchronous collaboration.
@@
To measure outcomes in synchronous vs. asynchronous collaboration, we propose a comprehensive evaluation strategy that looks at different output qualities. Namely, we are interested in measuring the consistency of outputs between synchronous and asynchronous coding, as well as the cohesiveness, and correctness of the resulting themes within each collaboration setting. Finally, we also perform a small-scale user study to draw insights from the experience of the coders under each collaboration setting.
In summary, we make two main contributions:
  1. 1.
    We demonstrate how the collaboration setting in which qualitative research takes place could affect the outcome of solutions proposed in the NLP literature, by testing two NLP-assisted inductive coding tools in synchronous and asynchronous coding contexts.
  2. 2.
    We propose a comprehensive evaluation strategy that measures different aspects of the outcome of NLP-assisted inductive coding tools, offering insights into the types and quality of themes that they are able to capture.
We hope that our findings will inform the future development of language technologies that are in line with the expectations and workflow of qualitative researchers, as well offer insights into how the next generation of qualitative research tools should be evaluated so that researchers are able to choose the right tool for their coding requirements.
@@ 2 Related Work
Our work builds on previous literature on computational solutions for qualitative coding. The overarching goal of these systems is to partially automate the qualitative coding process by either learning user-defined themes interactively Pacheco et al. (2023); Gao et al. (2023) or by inducing topics or clusters in an interactive, semi-supervised manner Fang et al. (2023); Smith et al. (2018). A separate but related line of work exemplified by Gao et al. (2024) uses language models to generate label recommendations as the user performs the coding process. While this system is explicitly designed for asynchronous collaboration, the systems we study differ in their ability to classify large portions of the dataset without requiring extensive supervision.
Our research addresses a real-world use case for qualitative researchers using HiTL systems and is informed by literature from the field of Human-Computer Interaction (HCI)  Jiang et al. (2021); Feuston and Brubaker (2021); Chen et al. (2018). HCI research centers the user as the focus of study in order to identify where and how best to place NLP tools in interactive systems. We build on these findings by proposing a comprehensive method for evaluating a given system for qualitative coding in addition to conducting user interviews. To this end, we also center part of our experiment on the user to uncover differences in the interactive experience between synchronous and asynchronous environments.
A common finding in the HCI studies surveyed is the researchers’ focus on ambiguity. In traditional settings, researchers can independently code a significant portion of the dataset and meaningfully identify areas of ambiguity. However, when working with large datasets and relying on NLP systems, researchers cannot easily uncover all such areas. Solutions have been proposed to either visualize codes (Drouhard et al., 2017) or rank document disagreement (Zade et al., 2018) regardless of dataset size. Our evaluation differs from these works since we look at different output qualities of the resulting themes and their assignments, combining signal from group overlaps, relationships in the semantic embedding space, as well as manual post-hoc evaluations. This method of evaluation highlights areas where the coders diverge both with each other and with the model, providing another perspective on the ambiguity question.
The motivating factor for our study is the popularization of HitL systems for qualitative coding in the NLP field. With the growing number of new interactive systems comes a need for standardized evaluation across different real-world environments. Previous evaluation methods introduced with new systems were generally ad hoc, with experiments conducted variously in group settings  Choo et al. (2013); Hoque and Carenini (2016); Smith et al. (2018), on individual participants  Rietz et al. (2020), and through platforms such as MTurk  Zade et al. (2018). Our contribution provides a standardized framework for performing experiments in synchronous and asynchronous settings, as well as a set of metrics for evaluating consistency, cohesiveness, and correctness in experimental results.
@@ 3 Interactive Systems
We evaluate two interactive theme discovery systems that use different approaches to accomplish the same task. This way, we ensure that our analysis can account for variation in NLP techniques and annotation interfaces.
3.1 Fang et al. (2023)
This system is a HitL, query-driven topic model (QDTM) (Fang et al., 2021) that allow researchers to iteratively adjust clusters using a set of refinement operations. Broadly, the QDTM is a semi-supervised topic modeling algorithm that accepts input queries to extract relevant concept words. The extracted concept words are used in a two-phase framework based off the Hierarchical Dirichlet Process (HDP) (Teh et al., 2004) to perform topic modeling. In the first phase, the HDP is used to infer one "parent topic" for each concept along with irrelevant topics, where each topic is defined by a set of weighted words. In the second phase, the model expands the parent topic into a set of subtopics using the Gibbs sampling process. The words for each subtopic are added to the parent topic to produce a final set of topics, which are used to assign a distribution of topics to each document.
Given an initial topic model, users can add, remove, or swap the order of words within each topic, which alters the probability of topic assignment for documents that contain that specified word. Similarly, users can split or merge topics based on topic words; doing so does not guarantee that assigned documents will be reassigned accordingly due to the stochastic nature of the model. Users can also remove documents from a given topic, but cannot directly add documents in the same manner. In each iteration, users may freely select refinement operations for each topic before applying all selections at once. Users can undo selections prior to applying refinements and previous models are also saved if users wish to return to a previous iteration.
The overarching goal of this system is to facilitate the discovery of an optimal topic model for a given input dataset. By applying a selection of refinements, users can compare the effects of different topic word allocations on the resulting model. Returning to previous models and applying new refinements creates a branching model, which provides users with a wide space for exploration. Once satisfied with a given iteration, the user can download that model with its document distribution.
3.2 Pacheco et al. (2023)
This system uses a two-stage framework to assist researchers in characterizing large textual collections. In the first stage, the system automatically proposes an initial partition of the data, such that textual instances that are thematically similar are clustered together. Users are then provided with a set of exploratory operations that allows them to evaluate the quality of the discovered partitions, as well as to further explore and partition the space by inspecting individual examples, finding similar instances, and using open text queries. Next, users are prompted to group identified patterns into general themes, and instantiate them by giving them a name. Once a theme is created, users are provided with a set of intervention operations to explain the themes using natural language, select good example instances, write down additional examples, and input or correct supporting concepts to characterize the theme assignments. The full set of operations supported by the system are outlined in App. A.
In the second stage, the system finds a mapping between the full set of instances and the themes instantiated by the users. It takes the information contributed by the users in the form of examples and supporting concepts, and learns to map instances to themes using a neuro-symbolic procedure implemented with DRaiL Pacheco and Goldwasser (2021). The main goal of this process is to condition new theme assignments not only on the embedding distance between instances and provided examples, but also leverage the additional judgements provided by users via supporting concepts. For example, when analyzing a corpus about the Covid-19 vaccine, users could point out that 80% of the examples for theme Government Distrust have a clear anti-vaccine stance. The system then uses this information to introduce inductive bias into the mapping procedure. DRaiL uses weighted first-order logic rules to express decisions and dependencies between different decisions, for example $w:\mathtt{stance(anti\_vax)}\rightarrow\mathtt{theme(gov\_distrust)}$. For more details about this implementation, we refer the reader to the original paper.
During this process, the systems keeps instances unassigned to a theme if there is not a good enough match. Then, a consecutive partitioning step is done over all unassigned instances. Finally, users have the option to repeat this two-stage process until convergence.
4 Coding Protocol
In order to study the performance of the two systems discussed in Section 3 under different collaboration settings, we design a protocol to conduct qualitative coding in the synchronous and asynchronous qualitative coding paradigms. We use the outcome of our coding described here to gather results, and then conduct an evaluation in the subsequent sections.
4.1 Participants
We recruited a group of 15 researchers in NLP and Computational Social Science, 3 female and 12 male, within the ages of 20 and 45. This group included professors at different levels of seniority, postdoctoral researchers, and graduate and undergraduate students from two different universities. All participants were either well versed in qualitative data analysis, or were explicitly trained by more senior researchers to perform the task.
4.2 Dataset
@@
We use a common dataset to perform qualitative coding using the Pacheco et al. (2023) and Fang et al. (2023) systems. This dataset was originally released by Pacheco et al. (2022), and consists of approximately $85,000$ Tweets about the COVID-19 vaccine. All tweets in this corpus were posted by users located in the United States, are uniformly distributed between Jan. and Oct. 2021, and contain predictions for vaccination stance (e.g. pro-vax, anti-vax) and morality frames (e.g. fairness/cheating and their actor/targets.) Haidt and Graham (2007); Roy et al. (2021).
4.3 Qualitative Coding Protocol
In order to compare the consistency between different collaborative settings, we need at least two sets of synchronous/asynchronous annotations per annotation system. To ensure compatibility between synchronous and asynchronous outcomes, we use the same starting clusters for each system. The hyperparameters used in the Fang et al. (2023) system are the same as in their study and we did not use any input queries prior to initialization. We scheduled annotations for each setup separately, with details outlined in the sections below.
4.3.1 Synchronous Coding
Coding for the Pacheco et al. (2023) system was done over Zoom, where participants met online and collaboratively discussed and annotated the dataset. Six participants were bifurcated into two groups of three individuals, with each group communicating only with their members during the annotation process. Coders followed the protocol outlined in the original paper, performing two consecutive rounds of interaction, each round lasting a total of 1 hour.
For the Fang et al. (2023) system, we selected a different set of six annotators and divided them into two separate groups of three individuals each. Annotators met with their group and collaboratively coded the data. Prior to the annotation session, participants were briefed about the different features of the system and how to use them. The annotators went over the pre-assigned topics proposed by the system, and modified them to make them more accurate and cohesive. Both groups of annotators reached saturation after multiple passes over the data, in sessions lasting around 2.5 hours.
4.3.2 Asynchronous Coding
The asynchronous coding processed followed the same outline as the synchronous coding for each system, except that it was conducted by three different annotators, and it was done individually without any collaboration. For each system, the annotators first met in-person and started annotations separately. The purpose of the initial co-located coding session was to troubleshoot any problems and solve any doubts that the participants may have. The participants then completed the annotation completely asynchronously and remotely.
Pacheco et al. (2023) Fang et al. (2023)
Jaccard Sim. Centroid Sim. Group Avg. Sim. Jaccard Sim. Centroid Sim. Group Avg. Sim.
Sync $\bf 0.36(0.19)$ $\bf 0.98(0.01)^{*}$ $\bf 0.52(0.07)^{*}$ $\bf 0.56(0.23)$ $\bf 0.98(0.05)^{**}$ $\bf 0.52(0.10)$
Asyn $0.30(0.22)$ $0.94(0.07)^{*}$ $0.44(0.10)^{*}$ $0.30(0.17)$ $0.96(0.05)^{**}$ $0.51(0.09)$

### v2

# Effects of Collaboration on the Performance of Interactive Theme Discovery Systems

@@ Abstract
NLP-assisted solutions have gained considerable traction to support qualitative data analysis. However, no unified evaluation framework exists which can account for the many different settings in which qualitative researchers may employ them. In this paper, we propose an evaluation framework to study the way collaboration settings may produce different outcomes across a variety of interactive systems. Specifically, we study the impact of synchronous vs. asynchronous collaboration using three different NLP-assisted qualitative research tools and present a comprehensive analysis of significant differences in the consistency, cohesiveness, and correctness of their outputs.
@@ 1 Introduction
Making sense of large amounts of textual data is a common task in many academic disciplines and is generally achieved through qualitative methods such as Thematic Analysis (Braun and Clarke, 2006) and Grounded Theory (Glaser et al., 1968). In these methods, researchers manually search through documents to identify common abstract themes that emerge from the data in a process commonly referred to as "inductive coding". However, manual inductive coding becomes infeasible as datasets increase exponentially in size, leading researchers to adopt Natural Language Processing (NLP) solutions to automate parts of the coding process Brady (2019); Hilbert et al. (2019).
@@
In the last decade, several NLP solutions have been proposed to support qualitative research methods. These systems offer different ways to systematically wrangle large amounts of textual data to identify common patterns or themes, such as discovering hidden semantic structures using topic models Smith et al. (2018); Fang et al. (2023), grouping documents based on similarity and broadcasting small sets of human annotations to the entire dataset Pacheco et al. (2023); Chew et al. (2023), or providing live recommendations to users as they code the data Dai et al. (2023); Gao et al. (2024). To ensure the agency of the researchers, each system seeks to strike a balance between automation and manual coding by incorporating some type of human-in-the-loop (HitL) strategy.
Figure 1: In this study, we measure the quality of coded themes using different interactive systems under different coding configurations.
@@
Generally, tools are evaluated to highlight the strengths and weaknesses of the proposed approach. Examples include evaluating topic coherence with and without human input Fang et al. (2023) and contrasting machine-assisted code-book generation with a manual code-book generation Dai et al. (2023). In this paper, we argue that these evaluation strategies are often too narrow and do not account for the different real-world settings in which qualitative researchers employ HitL systems. In particular, we seek to answer the following questions: (1) Does collaboration setting measurably affect the quality of resulting code-books? (2) Do these findings hold across different NLP approaches?
@@
Collaborative Qualitative Analysis (CQA) is a process where researchers work in teams to analyze a dataset to produce a consensus interpretation. Doing so can produce more reliable results, a richer understanding of the data, and insight that combines multiple perspectives Flick (2014). While there are several frameworks for conducting CQA, in this work we look at two common but contrasting collaboration settings. The first setting is defined by individual coders working separately before jointly consolidating their findings (asynchronous coding). The second setting is defined by coders working in groups and identifying themes through synchronous discussion (synchronous coding). Each setting has different strengths: asynchronous coding is more flexible as teams can work across different regions and time zones, whereas synchronous coding combines diverse perspectives and streamlines the annotation processes.
@@
We test these settings on three different NLP tools for interactive inductive coding by performing experiments in the synchronous and asynchronous contexts. We choose tools that use different underlying NLP methods for this work: (1) the Fang et al. (2023) system, which uses a human-in-the-loop topic modeling solution, (2) the Pacheco et al. (2023) system, which models emerging themes with respect to their relationship to other high-level concepts which can be defined by users, and (3) the Chew et al. (2023) system, which leverages large language models (LLMs) to judge the adequacy of code definitions, as well as to broadcast them to a large dataset. By way of testing such varied approaches, we achieve a more representative understanding of how different tools perform in different collaboration settings.
@@
To measure outcomes in synchronous vs. asynchronous collaboration, we propose a comprehensive evaluation strategy that looks at different output qualities. Namely, we are interested in measuring the consistency of outputs between synchronous and asynchronous coding, as well as the cohesiveness, and correctness of the resulting themes within each collaboration setting. Finally, we perform a small-scale user study to draw insights from the experience of the coders under each collaboration setting and make design recommendations according to our findings.
In summary, we make two main contributions: (1) We demonstrate how the collaboration setting in which qualitative research takes place affects the outcome of solutions proposed in the NLP literature, by testing three NLP-assisted inductive coding tools in synchronous and asynchronous coding contexts. (2) We propose a comprehensive evaluation strategy that measures different aspects of the outcome of NLP-assisted inductive coding tools, offering insights into the types and quality of themes that they are able to capture. We hope that our findings will inform the development of language technologies that are in line with the qualitative researcher workflows and offer insights for producing useful qualitative research tool evaluations.
@@ 2 Related Work
The overarching goal of the systems we investigate is to partially automate the qualitative coding process either by inducing topics in an interactive, semi-supervised manner Fang et al. (2023); Smith et al. (2018), by learning user-defined themes interactively Pacheco et al. (2023); Gao et al. (2023), or by prompting LLMs with natural language definitions of the observed themes Chew et al. (2023); Dai et al. (2023). A separate but related line of work exemplified by Gao et al. (2024) uses LLMs to generate label recommendations as users perform the coding process. While this system is explicitly designed for asynchronous collaboration, the systems we study differ in their ability to annotate large portions of the dataset without extensive supervision.
Our research addresses a real-world use case for qualitative researchers using HiTL systems and is informed by the Human-Computer Interaction (HCI) literature Jiang et al. (2021); Feuston and Brubaker (2021); Chen et al. (2018). A common finding in the HCI studies on computer-assisted qualitative research is the researchers’ focus on ambiguity. In traditional settings, researchers can independently code a significant portion of the dataset and meaningfully identify areas of ambiguity. However, when working with large datasets and relying on NLP systems, researchers cannot easily uncover all such areas. Solutions have been proposed to either visualize codes (Drouhard et al., 2017) or rank document disagreement (Zade et al., 2018) regardless of dataset size. Our evaluation differs from these works since we look at different output qualities of the resulting themes and their assignments, combining signal from group overlaps, relationships in the semantic embedding space, as well as manual post-hoc evaluations. This method of evaluation highlights areas where the coders diverge both with each other and with the model, providing another perspective on the ambiguity question.
The motivating factor for our study is the popularization of HitL systems for qualitative coding in the NLP field. With the growing number of new interactive systems comes a need for standardized evaluation across different real-world environments. Previous evaluation methods introduced with new systems have generally been ad hoc, with experiments conducted variously in group settings  Choo et al. (2013); Hoque and Carenini (2016); Smith et al. (2018), on individual participants  Rietz et al. (2020), and through platforms such as MTurk  Zade et al. (2018). Our contribution provides a standardized framework for performing experiments in synchronous and asynchronous settings, as well as a set of metrics for evaluating consistency, cohesiveness, and correctness in experimental results.
@@ 3 Interactive Systems
We identify three categories of NLP techniques used in interactive systems for qualitative coding with large datasets: topic models, relational approaches, and LLMs. These techniques may be applied in a variety of ways in interactive systems, but we put special focus on their ability to help code large datasets. To maximize coverage across systems, we select a representative system from each category to use in our experiments. In this section, we briefly describe the unique aspects of each category and introduce the selected system.
Topic Models. This category includes systems that use some variation of topic modeling to find emerging themes and facilitate document assignment. These systems benefit from the relative speed of the topic model, which allow users to quickly visualize and explore the dataset. Early exploration incorporated visualizations to help users adjust parameters (Chuang and McFarland, 2013), while later works implemented refinement operations that allow users to directly edit topic words and remove documents (Smith et al., 2018). However, topic modeling systems are limited by their lack of malleability and predictability. Refinement operations mostly edit topic words, which can have limited impact in the final results.
We select the HitL query-driven topic model (QDTM) system introduced by Fang et al. (2023). The topic model is initialized by providing input queries (i.e., words that represent concepts of interest for the user) which the model uses to generate the initial topics. In our experiments, users begin by iterating through each topic and naming them based on identified themes. Users then use a set of refinement operations to edit the topic model. They can merge and split topics based on topic words, add, remove, or reorder topic words, and remove documents from topics. The next iteration of the model is only produced when the users choose to apply refinements and the prior model is saved, allowing users to return to prior iterations to test different operations. Once satisfied with the state of the topic model, the user downloads the document distribution for that iteration. Details of our experimental settings can be found in App. A.
Relational Approaches. Relational approaches combine vector semantics and structured inference to model relationships between high-level concepts. Instead of treating themes as distributions over words (as topic models do), these frameworks define themes as distributions over generalized concepts. This reflects the inductive coding process, where researchers identify patterns and concepts that are then synthesized into more abstract themes. However, their computational complexity grows with the number of dependencies considered, which hinders their ability to quickly adapt during coding sessions. Further, they rely on users to define informative concepts, making them less suited for inexperienced researchers.
We select the system introduced by Pacheco et al. (2023), which uses a two-stage relational framework. In the first stage, the system automatically partitions the dataset based on semantic similarity. The users explore each partition to identify themes, assign "good" and "bad" example documents for each theme, and input or correct supporting concepts for each example. In the second stage, the system uses the provided examples and concept relations to map the remaining dataset, only leaving documents unmapped if no theme is a sufficiently good match. The assignment procedure follows as a structured inference approach, where dependencies between concepts and themes are explicitly modeled. The unmapped documents are repartitioned as in the first stage and users are prompted to review unmapped partitions again. The process is iterated until all documents are mapped. More details about the framework and our experimental configuration can be found in App. B
Large Language Models. LLMs are ideal candidates for interactive systems, especially for tasks such as qualitative coding where the model can be prompted to produce themes or explanations without ad-hoc training (Kojima et al., 2024). They have been used for theme recommendation (Gao et al., 2023), for code conflict resolution (Gao et al., 2024), and for automated document assignment (Xiao et al., 2023). However, the flexibility of LLM outputs also leads to hallucinations, which are only partially addressed by prompt engineering. Models further suffer from biases in training which are difficult to identify and impact their ability to produce quality labels or recommendations (Chen et al., 2018). Additionally, their massive size is prohibitive when working with large datasets due to the high cost of inference.
We select the framework introduced by Chew et al. (2023). In their protocol, the human coder first manually codes a representative subset of the data and drafts definitions for each code. The LLM is then prompted to label the data sample with the provided definitions. Agreement is calculated between human and model annotations using GWET’s $AC_{1}$ (Gwet, 2008). The prompt is then tweaked iteratively to achieve a satisfactory level of agreement, and the best-performing version is used to prompt the model to code the rest of the dataset. Details about our experimental settings can be found in App. D.
4 Study Design
To study the effects of different collaboration settings on the performance of the three selected systems, we design a protocol that can be used for both synchronous and asynchronous settings. For each system, we conduct three asynchronous experiments with one coder each and two synchronous experiments with three coders each for a total of 15 experiments. Evaluation metrics are calculated by comparing the resulting code-books within each experimental setting (e.g. the two code-books independently created by the two synchronous groups using the topic model). The rest of this section lays out the dataset, participant demographics and experimental protocol.
@@ 4 Study Design
Dataset. We use a common dataset for all experiments. It was originally released by Pacheco et al. (2022), and it consists of approximately $85,000$ Tweets about COVID-19 vaccines. All tweets in this corpus were posted by users located in the United States, are uniformly distributed between Jan. and Oct. 2021, and contain labels for vaccination stance (e.g. pro-vax, anti-vax) and morality frames (e.g. fairness/cheating and their actor/targets.) Roy et al. (2021).
Participants. We recruited a group of 20 researchers in NLP and Computational Social Science, 6 female and 14 male, between the ages of 20 and 45. This group included professors at different levels of seniority, postdoctoral researchers, and graduate and undergraduate students from two different universities. This group covers the range of researchers likely to use interactive coding systems. All participants were either well-versed in qualitative data analysis, or were explicitly trained by senior researchers to perform the task.
Coding Protocol. At the start of each experiment, participants were provided with a demonstration of all the operations in their respective systems. Every system starts with an initial partition of the data, so participants were instructed to read the first 25 samples in each partition, and manually create/name any themes they identified before freely exploring the rest of the dataset and start performing operations to find more themes.
In the topic model experiments, we suggested that participants merge and split topics based on their identified themes before making fine-grained refinements. They were then asked to refine the topic model based on their identified themes such that every topic corresponds to a unique theme. They kept re-running the model and making refinements until they were satisfied with the results, or until they failed to effect any meaningful changes.
In the relational system experiments, participants were tasked with selecting example documents for each identified theme, as well as determining concept relations for them. Following Pacheco et al. (2023), the supporting concepts considered were vaccination stances and morality frames (e.g., the identified theme “natural immunity” has an “anti-vax” stance, and is tied to the “purity” frame). Once participants were satisfied with their themes and selections, the system automatically coded the rest of the dataset. Unmapped examples were repartitioned and returned to the participants for a second (and last) round of coding.
In the experiments for the LLM-based system, participants produced natural-language definitions for each identified theme and selected a set of good examples for them. We then prompted the LLM with different task-prompt templates to find the best prompt for each set of participant-generated definitions, which was then used to code the rest of the dataset. Details of the templates, as well as the human-model agreement for the best template can be found in Appendix D.
Topic Model Relational LLM-Based
Jaccard Centroid Group Avg. Jaccard Centroid Group Avg. Jaccard Centroid Group Avg.
Sync $\bf 0.56(0.23)$ $\bf 0.98(0.05)^{**}$ $\bf 0.52(0.10)$ $\bf 0.36(0.19)$ $\bf 0.98(0.01)^{*}$ $\bf 0.52(0.07)^{*}$ $0.14(0.08)$ $0.98(0.03)$ $0.44(0.03)$
Async $0.30(0.17)$ $0.96(0.05)^{**}$ $0.51(0.09)$ $0.30(0.22)$ $0.94(0.07)^{*}$ $0.44(0.10)^{*}$ $\bf 0.17(0.11)$ $0.98(0.02)$ $\bf 0.45(0.03)$